# Extended Physics Informed Neural Network for Hyperbolic Two-Phase Flow in Porous Media

### v1

@@ Abstract
The accurate solution of nonlinear hyperbolic partial differential equations (PDEs) remains a central challenge in computational science due to the presence of steep gradients, discontinuities, and multiscale structures that make conventional discretization-based solvers computationally demanding. Physics-Informed Neural Networks (PINNs) embed the governing equations into the learning process, enabling mesh-free solution of PDEs, yet they often struggle to capture steep gradients, discontinuities, and complex nonlinear wave interactions. To address these limitations, this study employs the Extended Physics-Informed Neural Network (XPINN) framework to solve the nonlinear Buckley-Leverett equation with a nonconvex flux function, which models immiscible two-phase flow in porous media. The computational domain is dynamically decomposed in space and time into evolving pre-shock and post-shock regions, allowing localized subnetworks to efficiently learn distinct flow behaviors. Coupling between subnetworks is achieved through the Rankine-Hugoniot jump condition, which enforces physically consistent flux continuity across the moving shock interface. Numerical experiments demonstrate that the proposed XPINN approach accurately captures discontinuous saturation fronts and compound wave interactions without requiring artificial diffusion or entropy corrections. Compared to standard PINNs, the XPINN framework achieves superior stability, faster convergence, and enhanced resolution of nonlinear wave dynamics using smaller, domain-specific models with fewer trainable parameters, establishing it as an effective and scalable tool for solving challenging hyperbolic PDEs in multiphase flow problems. The code of this work is available on github.com/saifkhanengr/XPINN-for-Buckley-Leverett.
@@ 1 Introduction
In this paper, we aim to apply XPINNs to the Riemann problem characterized by nonconvex flux function with multiple inflection points. The Buckley-Leverett equation provides an ideal test case, as it is a nonlinear hyperbolic PDE modeling multiphase flow in porous media that generates steep saturation fronts and shock-like structures [23, 49]. Such features are notoriously difficult for standard PINNs to capture, as highlighted by [43, 44, 45, 46, 47, 48, 49]. We employ XPINNs to study the Buckley-Leverett problem, demonstrating how the combination of physics-informed learning and subdomain decomposition enables efficient training and accurate resolution of discontinuous saturation fronts.
The remainder of this paper is organized as follows. In Section 2, we introduce the XPINN framework and highlight how it differs from standard PINNs in solving deterministic PDEs. Section 3 provides an overview of the two-phase transport model along with the governing hyperbolic PDE. The results for the transport problem are presented in Section 4. Finally, Section 5 summarizes our findings and offers a discussion of the outcomes.
2 Methodoloyy
@@ 2.2.1 Loss Functions and Interface Conditions for XPINNs
$J(\theta_{q})$ $=\underbrace{\lambda_{u_{q}}L_{u_{q}}(\theta_{q})}_{\text{Data}}+\underbrace{\lambda_{F_{q}}L_{F_{q}}(\theta_{q})}_{\text{Residual}}+\underbrace{\lambda_{IF_{q}}L_{R}(\theta_{q})}_{\text{Interface Residual}}+\underbrace{\lambda_{I_{q}}L_{\text{avg}}(\theta_{q})}_{\text{Interface Average (Optional)}}.$ (11)
@@ 4 Computational Experiments
This section presents the computational setup and numerical experiments performed using the proposed XPINN framework. The domain is dynamically decomposed in both space and time into two subdomains—pre-shock and post-shock regions—to accurately capture the evolving discontinuity and distinct flow characteristics across the shock front.
The model is implemented for the one-dimensional Buckley-Leverett problem over $x\in[0,1]$ and $t\in[0,1]$ with mobility ratio $M=2$. The dataset includes 200 points for the initial condition at $t=0$, 200 for the boundary at $x=0$, 2000 collocation points each for the pre- and post-shock regions, and 200 points along the shock interface connecting them, as shown in Figure 2. Moreover, the computational domain is decomposed dynamically so that the subdomains evolve with the moving shock. For each spatial and temporal instance, a single collocation point along the shock interface is maintained, while multiple points are assigned to the pre-shock and post-shock regions. As the shock propagates, the collocation points are updated accordingly, ensuring that the data distribution adapts to the changing flow structure, as illustrated in Figure 3. Since the interface in this case corresponds to the shock region (a discontinuity), the Interface Average term of Eq. 11 is not enforced, consistent with the reasoning provided in Section 2.2.1.
Figure 2: Training Data Distribution.
@@ 5 Conclusion and Future Work
This work applied the Extended Physics-Informed Neural Network (XPINN) framework to the Buckley-Leverett equation with a nonconvex flux, demonstrating that dynamic domain decomposition in space-time—partitioning the domain into evolving pre-shock and post-shock regions—enables each subnetwork to learn localized solution structure and collectively represent sharp saturation fronts. Enforcing the Rankine-Hugoniot jump condition along the moving interface provided a physically consistent coupling between subnetworks and allowed the model to capture discontinuous behavior without introducing artificial diffusion or entropy constraints.

### v2

@@ Abstract
The accurate solution of nonlinear hyperbolic partial differential equations (PDEs) remains challenging due to steep gradients, discontinuities, and multiscale structures that make conventional solvers computationally demanding. Physics-Informed Neural Networks (PINNs) embed the governing equations into the learning process, enabling mesh-free solution of PDEs, yet they often struggle to capture steep gradients, discontinuities, and complex nonlinear wave interactions. To address these limitations, we employ the Extended Physics-Informed Neural Network (XPINN) framework to solve the nonlinear Buckley-Leverett equation with a nonconvex flux, modeling immiscible two-phase flow in porous media. The computational domain is dynamically decomposed in space and time into evolving pre-shock and post-shock subdomains, allowing localized subnetworks to efficiently learn distinct flow behaviors, with coupling enforced via the Rankine-Hugoniot jump condition to ensure physically consistent flux continuity. We compare XPINN with standard PINNs and its variants, including PINN with artificial viscosity, PINN with Welge construction, and PINN with the Oleinik entropy condition, and across all cases, XPINN consistently outperforms the other methods, accurately resolving sharp fronts and capturing the correct physical behavior. Importantly, XPINN achieves this using the simpler Adam optimizer, whereas some PINN variants require more complex or higher-order strategies such as L-BFGS to reach comparable accuracy, demonstrating that XPINN is a robust and scalable approach for challenging hyperbolic PDEs without artificial diffusion or entropy corrections. The code is available at github.com/saifkhanengr/XPINN-for-Buckley-Leverett.
@@ 1 Introduction
In this paper, we apply XPINNs to the Riemann problem characterized by a nonconvex flux function with multiple inflection points. The Buckley-Leverett equation provides an ideal test case, as it is a nonlinear hyperbolic PDE modeling multiphase flow in porous media that generates steep saturation fronts and shock-like structures [23, 49]. Such features are notoriously difficult for standard PINNs to capture, often requiring artificial diffusion, entropy corrections, or complex network architectures, and higher-order optimizers such as L-BFGS [43, 44, 45, 46, 47, 48, 49]. We employ XPINNs to study the Buckley-Leverett problem, demonstrating how the combination of physics-informed learning and dynamic subdomain decomposition enables efficient training, accurate resolution of discontinuous saturation fronts, and superior performance compared to standard PINNs, even when using the simpler Adam optimizer.
The remainder of this paper is organized as follows. In Section 2, we introduce the XPINN framework and highlight how it differs from standard PINNs in solving deterministic PDEs. Section 3 provides an overview of the two-phase transport model along with the governing hyperbolic PDE. The results for the transport problem, including a comparison of XPINN with standard PINNs and its variants, are presented in Section 4. Finally, Section 5 summarizes our findings and discusses the outcomes.
2 Methodoloy
@@ 2.2.1 Loss Functions and Interface Conditions for XPINNs
$J(\theta_{q})$ $=\underbrace{\lambda_{u_{q}}L_{u_{q}}(\theta_{q})}_{\text{Data}}+\underbrace{\lambda_{F_{q}}L_{F_{q}}(\theta_{q})}_{\text{Residual}}+\underbrace{\lambda_{IF_{q}}L_{R}(\theta_{q})}_{\text{Interface Residual}}+\underbrace{\lambda_{I_{q}}L_{\text{avg}}(\theta_{q})}_{\text{Interface Average (Optional)}}$ (11)
[…]
@@ 4 Computational Experiments
This section presents the computational setup and numerical experiments performed using the proposed XPINN framework, followed by a comparison against standard PINN and its variants, including PINN with artificial viscosity, PINN with Welge construction, and PINN with the Oleinik entropy condition.
4.1 Results and Analysis of XPINNs
4.1.1 XPINN With Interface Residual
Here, we show the results of XPINN obtained using Equation 11. The domain is dynamically decomposed in both space and time into two subdomains—pre-shock and post-shock regions—to accurately capture the evolving discontinuity and distinct flow characteristics across the shock front. The model is implemented for the one-dimensional Buckley-Leverett problem over $x\in[0,1]$ and $t\in[0,1]$ with mobility ratio $M=2$. The dataset includes 200 points for the initial condition at $t=0$, 200 for the boundary at $x=0$, 2000 collocation points each for the pre- and post-shock regions, and 200 points along the shock interface connecting them, as shown in Figure 2. Moreover, the computational domain is decomposed dynamically so that the subdomains evolve with the moving shock. For each spatial and temporal instance, a single collocation point along the shock interface is maintained, while multiple points are assigned to the pre-shock and post-shock regions. As the shock propagates, the collocation points are updated accordingly, ensuring that the data distribution adapts to the changing flow structure, as illustrated in Figure 3. Since the interface in this case corresponds to the shock region (a discontinuity), the Interface Average term of Eq. 11 is not enforced, consistent with the reasoning provided in Section 2.2.1.
Figure 2: Training Data Distribution for XPINN.
@@ 4.1.1 XPINN With Interface Residual
4.1.2 XPINN Without Interface Residual
Here, we present the results obtained using Equation 11 without enforcing the interface residual. XPINN without the interface residual is similar to a standard PINN, with the main difference being that two separate PINNs are used—one for each subdomain—instead of a single PINN for the entire domain. All other settings are the same as those described in Section 4.1.1 (with interface). Since the interface corresponds to a discontinuity, the average interface term is not applied as explained earlier. Figure 7 shows the results for architecture 2 with the hyperparameters summarized earlier in Table 2.
Figure 7: Model results for Architecture 2 without Residual Interface term.
As seen in Figure 7, XPINN without residual interface accurately captures the shock location without the residual interface term, though post-shock saturation does not reach zero. We experimented with different values of $M$ (from 0.5 to 30) and observed similar results. So, we deem that the use of separate subnetworks on each subdomain introduces enough representational freedom for a steep interface to form, allowing the model to approximate the correct shock location even without the Rankine-Hugoniot constraint. However, this accuracy is limited: without the residual interface, the interface states are not physically anchored, and the post-shock value in subnetwork 2 remains slightly above zero. This occurs because the solution starts from the interface value, and the PDE residual in the flat post-shock region cannot push it toward the exact physical value: any constant in this region satisfies the PDE locally, so the residual is near zero and the optimizer receives almost no gradient to correct the output. Table 3 provides a detailed mathematical illustration. Therefore, the residual interface term is essential not for locating the shock but for enforcing the correct left and right states across it, which sets the post-shock value to the physically accurate zero-saturation flat region.
Table 3: Post-shock region ($u=C$, a constant = 0) comparison without and with interface residual
Component Without Interface Residual With Interface Residual
Field Solution $u(x,t)=C$ $u(x,t)=C$
Spatial Derivative $\dfrac{\partial u}{\partial x}=0$ $\dfrac{\partial u}{\partial x}=0$
Time Derivative $\dfrac{\partial u}{\partial t}=0$ $\dfrac{\partial u}{\partial t}=0$
Flux Derivative $\dfrac{\partial f(u)}{\partial x}=f^{\prime}(C)\,\dfrac{\partial u}{\partial x}=0$ $\dfrac{\partial f(u)}{\partial x}=f^{\prime}(C)\,\dfrac{\partial u}{\partial x}=0$
PDE Residual $R_{\text{PDE}}=\dfrac{\partial u}{\partial t}+\dfrac{\partial f(u)}{\partial x}=0$ $R_{\text{PDE}}=\dfrac{\partial u}{\partial t}+\dfrac{\partial f(u)}{\partial x}=0$
Interface Residual $R_{\text{interface}}=0$ $R_{\text{interface}}=(u_{\text{post-shock}}-u_{\text{interface}})\neq 0$
Total Residual $R_{\text{total}}=0$ $R_{\text{total}}=R_{\text{PDE}}+R_{\text{interface}}\neq 0$
Interpretation The PDE is satisfied exactly even if $C\neq 0$. Optimizer gets almost no gradient to correct post-shock value. The interface residual creates a gradient at the shock interface, which propagates into the post-shock region and forces the network toward the correct value ($C=0$).
4.2 Comparision of XPINN with Standard PINN and its variants
In this section, we compare the results of the XPINN approach with the Standard PINN and its variants, including PINN with artificial viscosity (diffusivity term), PINN with Welge construction, and PINN with the Oleinik entropy condition. To ensure a fair comparison, the same dataset and training points used for XPINN (Section 4.1.1) are employed for all Standard PINN variants. Before presenting the results and associated mathematical formulations, we first illustrate the data distribution used for the Standard PINN. The setup is identical to that described in the XPINN section, except that here we consider a single unified dataset rather than separate subdomains, as shown in Figure 8.
Figure 8: Training Data Distribution for Standard PINN.
4.2.1 Standard PINN
Equation 11 shows the Standard PINN part which is
$J(\theta_{q})=\underbrace{L_{u_{q}}(\theta_{q})}_{\text{Data}}+\underbrace{L_{F_{q}}(\theta_{q})}_{\text{Residual}}$ (26)
where, $L_{u_{q}}$ and $L_{F_{q}}$ are the same as defined by Equation 12 and 13.
4.2.2 PINN with Diffusivity Term
A second-order artificial viscosity (diffusivity) term is incorporated into the residual [43]. Accordingly, the residual term in Equation 19 is modified as:
$L_{F_{q}}(\theta_{q})=\frac{1}{N_{F_{q}}}\sum_{i=1}^{N_{F_{q}}}\bigg|\frac{\partial S_{w_{\theta_{q}}}}{\partial t}+\frac{\partial f_{w}(S_{w_{\theta_{q}}})}{\partial x}-\epsilon\frac{\partial^{2}S_{w_{\theta_{q}}}}{\partial x^{2}}\bigg|^{2},$ (27)
where $\epsilon>0$ denotes the diffusion coefficient. Based on [43], $\epsilon$ is set to $2.5\times 10^{-3}$.
4.2.3 PINN with Welge Construction
Using the Welge construction to enforce the convex hull of the fractional flow [52], the modified fractional flow function is expressed as:
$\tilde{f}(S_{w})=\begin{cases}0,&S_{w}=0\\[2.84526pt] \dfrac{S_{w}}{f(S^{*})},&0<S_{w}\leq S^{*}\\[2.84526pt] f(S_{w}),&S_{w}>S^{*}\end{cases}$ (28)
4.2.4 PINN with Oleinik Entropy Condition
The Oleinik entropy condition is applied to maintain physically admissible shock solutions [45], leading to the following convex-hull representation of the fractional flow:
$\tilde{f}(S_{w})=\begin{cases}\sigma S_{w},&0\leq S_{w}<S^{*}\\[2.84526pt] f_{w}(S_{w}),&S_{w}\geq S^{*}\end{cases}$ (29)
where, $S^{*}$ denotes the saturation at the shock front, and $\sigma$ represents the corresponding shock speed, which is given by
$\sigma=\frac{f(S^{*})}{S^{*}}.$ (30)
Building upon these formulations, we conducted comprehensive experiments comparing all methods for the case of $M=2$. It is important to note that all models were configured with approximately the same number of trainable parameters and identical hyperparameters, as detailed in Table 4, ensuring a fair comparison across methods. The predictive results presented in Figure 9 and as reported in prior studies, the Standard PINN approach (Row 3) fails to accurately capture the solution and post shock saturation. Similarly, the PINN with artificial diffusivity (Row 4) demonstrates limited improvement and fails to resolve the sharp interface adequately. While [43] reported successful predictions using the L-BFGS optimizer, our comparative analysis employs the Adam optimizer consistently across all methods. The PINN variants incorporating Welge construction (Row 5) and Oleinik entropy condition (Row 6) yield physically plausible solutions with moderate accuracy, though their performance remains inferior to the XPINN method (Row 1), which demonstrates superior resolution of the shock front.
Table 4: Model configuration and hyperparameters for XPINN and PINN variants.
Component XPINN (2 sub-domains) Standard PINN PINN Diffusivity PINN Welge PINN Oleinik
# Hidden Layers 2 and 1 2 2 2 2
# Neurons (30, 20) and (10) 30, 22 30, 22 30, 22 30, 22
# Epochs 5,000 5,000 5,000 5,000 5,000
Total Trainable Parameters 777 798 798 798 798
Total Collocation Points 4,600 4,600 4,600 4,600 4,600
Learning Rate $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$ $10^{-3}$
Optimizer Adam Adam Adam Adam Adam
Adaptive Activation Function Tanh Tanh Tanh Tanh Tanh
Figure 9: Comparision of XPINN with Standard PINN and its variants for $M=2$.
The convergence behavior for $M=2$, illustrated in Figure 10, reveals that XPINN exhibits significantly smoother and more stable optimization dynamics compared to other methods, which display substantial oscillations during training. This enhanced convergence property is further substantiated by the quantitative metrics presented in Figure 11, which compares all methods using L1 and L2 error norms on test dataset alongside training time. Although XPINN requires marginally more computational time, it achieves the lowest error metrics among all compared methods, demonstrating an excellent trade-off between accuracy and computational cost.
Figure 10: Convergence behaviour of all methods for $M=2$.
Figure 11: Quantitative metrics of all methods for $M=2$.
Notably, for the more challenging case of $M=30$ (Figure 12), the performance gap becomes even more pronounced. While the Oleinik entropy condition and other PINN variants fail to produce accurate predictions under these more demanding conditions, the XPINN approach maintains robust performance, successfully resolving the sharper shock front and demonstrating the method’s scalability and robustness across different mobility ratio scenarios. The hyperparameters used in this study are the same as those shown in Table 4.
Figure 12: Comparision of XPINN with Standard PINN and its variants for $M=30$.
@@ 5 Conclusion and Future Work
This work applied the Extended Physics-Informed Neural Network (XPINN) framework to the Buckley-Leverett equation with a nonconvex flux, demonstrating that dynamic domain decomposition in space-time—partitioning the domain into evolving pre-shock and post-shock regions—enables each subnetwork to learn localized solution structure and collectively represent sharp saturation fronts. Enforcing the Rankine-Hugoniot jump condition along the moving interface provided a physically consistent coupling between subnetworks and allowed the model to capture accurate behavior without introducing artificial diffusion or entropy constraints.
We compared the results of the XPINN approach with the Standard PINN and its variants, including PINN with artificial viscosity (diffusivity term), PINN with Welge construction, and PINN with the Oleinik entropy condition. Across all cases and mobility ratios, XPINN consistently outperformed the other methods, accurately resolving sharp fronts and capturing the correct physical behavior, while the other PINN variants struggled under increasingly challenging conditions. Significantly, XPINN achieved this using the simpler Adam optimizer, whereas some PINN variants required more complex or higher-order optimization strategies, such as L-BFGS, to reach comparable accuracy.